# Integrating Multimodal Data for Joint Generative Modeling of Complex Dynamics

### v2

@@ 1 Introduction
In general, multimodal data integration is a thriving topic in many areas of artificial intelligence and ML research (Ahuja et al., 2017; Bhagwat et al., 2018; Baltrusaitis et al., 2019; Sutter et al., 2020; Shi et al., 2020; Antelmi et al., 2018; Liang et al., 2022; Lipkova et al., 2022; Xu et al., 2023; Steyaert et al., 2023; Warner et al., 2023; OpenAI, 2023). Generative models fusing multiple data channels produce a common latent code which allows for cross-modal prediction or output generation (e.g., images from language Amrani et al. (2021); Radford et al. (2021)), improves inference by complementing information too noisy or missing in one channel through recordings from other modalities (Qian et al., 2022), or may reveal interesting links among observed modalities (Liang et al., 2015).
@@
DSR, however, goes beyond mere forecasting in that we request a dynamical model of the data-generating process which captures the temporal and geometrical structure of the true underlying system (Koppe et al., 2019b; Abarbanel, 2022; Platt et al., 2023). A successful DS reconstruction can reveal important invariant topological and geometrical characteristics of the state space the DS lives in, its long-term temporal behavior (attractors), and its sensitivity to perturbations and parameter variations. It provides a mechanistic surrogate for the observed system which can be further analyzed, simulated, perturbed, or lesioned to obtain additional insight into its inner workings (Durstewitz et al., 2023). Often special training techniques (Mikhaeil et al., 2022; Brenner et al., 2022; Hess et al., 2023; Abarbanel, 2013, 2022; Vlachas & Koumoutsakos, 2023), regularization approaches (Schmidt et al., 2021; Platt et al., 2022; Jiang et al., 2023; Platt et al., 2023), or structural priors as in physics-informed neural networks (PINNs Raissi et al. (2019)), are required to achieve these goals. Plain ‘out-of-the-box’ gradient descent techniques or standard variational inference will often fail to reconstruct the underlying system’s long-term properties (Arribas et al., 2020; Brenner et al., 2022; Hess et al., 2023; Platt et al., 2023; Jiang et al., 2023; Durstewitz et al., 2023), as these standard procedures do not sufficiently constrain the vector field to prevent true and reconstructed trajectories from quickly diverging. This is especially a problem in chaotic systems, where it is related to the exploding or vanishing gradient problem in RNN training (Hochreiter & Schmidhuber, 1997; Hochreiter et al., 2001; Bengio et al., 1994; Mikhaeil et al., 2022; Hess et al., 2023). Control-theoretic techniques are often used to keep model-generated trajectories on track, or to synchronize them with those produced by the observed system, whilst training (Voss et al., 2004; Abarbanel et al., 2009; Abarbanel, 2013; Verzelli et al., 2021; Platt et al., 2021; Abarbanel, 2022). Similar ideas underlie the recently proposed methods of sparse (Mikhaeil et al., 2022; Brenner et al., 2022) and generalized (Hess et al., 2023) teacher forcing (STF and GTF, respectively) for learning chaotic DS, which balance model-generated and data-inferred states during training in a way that optimally controls gradient flows and trajectory divergence (Mikhaeil et al., 2022; Hess et al., 2023; Eisenmann et al., 2023).
@@
Although important in many areas of science, an efficient framework for multimodal data integration for the purpose of DS reconstruction is currently essentially lacking. A particular challenge here is to reconcile highly efficient training techniques for DS reconstruction like STF or GTF, which usually rest on an invertible decoder model, with the need to infer a common latent code from many different types of data. Here we address this challenge through a novel formulation of the multimodal data integration problem for DS reconstruction: Building on the success of variational autoencoders (VAEs) for multimodal integration (Baltrusaitis et al., 2019; Wu & Goodman, 2018; Sutter et al., 2020), we use multimodal VAEs (MVAEs) to construct a common latent representation from random variables following different distributional models. Crucially, however, rather than constructing a sequential VAE process that directly operates on this latent code (Kramer et al., 2022), here we employ the MVAE solely to create a multimodal teacher signal for guiding a DS reconstruction model in training. Special loss terms and shared decoder models ensure consistency between the MVAE’s and reconstruction model’s latent codes. This not only efficiently exploits the information available in multiple different data streams for DS reconstruction (as compared to a variety of alternative techniques), but for the first time, to our knowledge, enables DS reconstructions of chaotic systems from categorical information alone.
2 Multimodal Teacher Forcing (MTF) framework
Techniques like STF (Brenner et al., 2022; Mikhaeil et al., 2022) or GTF (Hess et al., 2023) enabled reconstructions even from challenging real-world data on which many previous methods failed. Hence, the key idea for making DS reconstruction from multiple, statistically distinct data sources work is the generation of a multimodal TF signal for efficient training of a reconstruction model. For this we utilize an MVAE which is then trained jointly with the DS reconstruction (DSR) model through a set of shared decoder models. This general framework, illustrated in Fig. 1, is very flexible and modular. In principle it can be used with any type of DSR model for approximating the flow of the observed system, as well as with any set of decoder and encoder models.
Figure 1: MTF setup. Multimodal observations are translated via an encoder into a common latent representation, which is used for sparse TF in the DSR model’s latent space. The latent trajectory is then mapped back into observation space via modality-specific decoder models.
@@ 2 Multimodal Teacher Forcing (MTF) framework
While completely flexible, for our specific experiments we chose the recently introduced ‘dendritic piecewise linear RNN (dendPLRNN)’ (Brenner et al., 2022) as a reconstruction model, as it is both expressive and has a mathematically tractable, piecewise linear design (Monfared & Durstewitz, 2020; Eisenmann et al., 2023) (other types of RNNs, such as LSTMs or GRUs, may be used and give similar results, see Suppl. Table B3). The dendPLRNN is defined by the $M$-dimensional latent process equation
@@
To infer this latent process equation jointly from multiple data modalities, the dendPLRNN is connected to different decoder models that take the distinct distributional properties of each modality into account. While our formulation is general and can work with any combination of continuous and/or discrete data types, for our exposition we consider as example time series of multivariate Gaussian ($\bm{X}$), ordinal ($\bm{O}$), [^1] Ordinal data, as often encountered in psychology or economics (Likert, 1932), are non-metric but still ordered. and count ($\bm{C}$) nature of length $T$, $\bm{Y}=\{\{\bm{x}_{1},\dots,\bm{x}_{T}\};\{\bm{o}_{1},\dots,\bm{o}_{T}\};\{\bm{c}_{1},\dots,\bm{c}_{T}\}\}$. In this case we may take as decoders a linear-Gaussian model, a cumulative link model for ordinal data, and a log-link function for Poisson data (see Methods A.1):
@@
$\bm{c}_{t}|\bm{z}_{t}$ $\sim\operatorname{Poisson}(\lambda(\bm{z}_{t})).$ (4)
@@
Training RNNs on time series is generally challenging due to the exploding and vanishing gradient problem (Bengio et al., 1994). This becomes particularly severe in DS reconstruction as for chaotic systems exploding gradients cannot be avoided even in principle, due to the positive maximum Lyapunov exponent resulting in exponential divergence of trajectories (Mikhaeil et al., 2022; Hess et al., 2023). Yet, at the same time batch length cannot be too short, as the training algorithm will then fail to capture the system’s long term behavior and limit sets (Mikhaeil et al., 2022; Platt et al., 2022, 2023). STF (Mikhaeil et al., 2022) (or GTF (Hess et al., 2023)) therefore balance loss and trajectory divergence with the need to capture relevant long time scales in an optimal way. They do so by replacing forwarded-iterated latent states $\bm{z}_{t}$ by data-inferred states $\bm{\hat{z}}_{t}$ at strategically chosen time points (determined from the empirically estimated maximum Lyapunov exponent in the case of STF (Mikhaeil et al., 2022), or by averaging forward-propagated and data-inferred states in an optimal way in the case of GTF (Hess et al., 2023)). Data-inferred states are obtained by inverting the decoder model (e.g., $\bm{\hat{z}}_{t}=(\bm{B}^{T}\bm{B})^{-1}\bm{B}^{T}\bm{x}_{t}$ for a simple linear-Gaussian model as in eq. 2, i.e. taking the pseudo-inverse). This forcing then is only applied during model training, not at run time (see Methods A.1 for further details).
Inverting the decoder model is in general not possible, however, in particular for discrete random variables. Moreover, in the case of multiple simultaneously observed data modalities, it is unclear how to combine the different data modalities to obtain an optimal estimate $\bm{\hat{z}}_{t}$. We therefore need another means to create a (sparse) TF signal if we would like to utilize STF or GTF for training. Here we achieve this through an MVAE for building a joint latent representation over the different data types, $\bm{\tilde{z}}_{t}\sim p(\bm{\tilde{z}}_{t}|\bm{X},\bm{O},\bm{C})$, which can be used as a TF signal at time $t$. We denote the MVAE states by $\bm{\tilde{z}}_{t}\in\mathbb{R}^{K}$ to avoid confusion with the latent dynamical process $\bm{z}_{t}\in\mathbb{R}^{M}$ generated by the reconstruction model. [^2] Note that in general we do not assume the MVAE and reconstruction model latent codes to have the same dimensionality, i.e. $K\leq M$, to accommodate different forms of TF, see Methods A.1 for further details To ensure consistency between the DSR model’s and the MVAE’s latent states, the MVAE is coupled to the observations through the very same set of decoder models
@@
sharing all decoder parameters with the DSR model (see Fig. 1). The MVAE is trained by minimizing the negative Evidence Lower Bound (ELBO)
@@
using the reparameterization trick for latent random variables (Kingma & Welling, 2014), where $\mathbb{H}_{q_{\bm{\phi}}}$ is the entropy of the approximate posterior, and assuming conditional independence of the observations given the latent states. We tested various choices for the encoder model ${q_{\bm{\phi}}}(\bm{\tilde{Z}}|\bm{Y})$, achieving best results with a CNN (see Table B2 for results and Methods A.1 for architectural details).
@@
We thus have specified ${p_{\bm{\phi}}}(\bm{Y}|\bm{\tilde{Z}})$ through our choice of decoder models, and parameterized ${q_{\bm{\phi}}}(\bm{\tilde{Z}}|\bm{Y})$ by a CNN. But what is the best choice for the prior $p_{\bm{\theta}}(\bm{\tilde{Z}})$? Crucially, to further enforce consistency between the MVAE and DSR model latent codes, $\bm{\tilde{Z}}$ and $\bm{Z}$, respectively, we assume that $p_{\bm{\theta}}(\bm{\tilde{Z}})$ is given through the DSR model (the dendPLRNN in our case). Specifically, in line with the Gaussian assumptions in the dendPLRNN, eq. 1, we assume for the second term in eq. 8
$\mathbb{E}_{q_{\bm{\phi}}}[\log p_{\bm{\theta}}(\bm{\tilde{Z}})]\approx$ $\frac{1}{L}\sum_{l=1}^{L}\sum_{t=1}^{T}-\frac{1}{2}\bigl{(}\log|\bm{\Sigma}|$ (9)
@@
The total training loss is now given by the MVAE’s negative ELBO, eq. 8, combined with the reconstruction loss (negative log-likelihood) of the DSR model (funneled through the same set of decoder models), i.e. by adding the term $-\log p_{\bm{\theta}}(\bm{Y}|\bm{Z})=:\mathcal{L}_{\text{DSR}}$ to eq. 8, where $\bm{Z}$ are the latent states of the DSR model:
Whilst training, we then use $\bm{\tilde{z}}_{t}$ as our TF signal for guiding the DSR model. Specifically, for STF we replace (a subset of the) latent states $\bm{z}_{t}$ by $\bm{\tilde{z}}_{t}$ at strategically chosen times $l\tau+1,\ l\in\mathbb{N}_{0}$, as illustrated in Fig. 1 (similarly, for GTF we may take a weighted average; see (Hess et al., 2023) for details). Note that our usage of an MVAE here thus profoundly differs from the more conventional use in, for instance, sequential VAEs (Sutter et al., 2020; Kramer et al., 2022). Ablation studies further confirm that all components of the loss (Fig. A3), as well as an optimal choice for the TF interval (Fig. A5), are crucial to the success of the approach. Specifically, optimal TF intervals were associated with smooth and well navigable loss landscapes (Fig. A4).
3 Results
3.1 Evaluating MTF: Comparison to other algorithmic strategies
We first compared MTF on moderately challenging multimodal DS reconstruction tasks [^3] Moderately challenging because many of the comparison methods actually broke down on the more difficult setups. to a variety of other possible algorithmic strategies that may be considered, all of which lacking one or more key ingredients of our methodology. For this, a training and a test set of $100,000$ time steps were created from a Lorenz-63 and a Rössler system with $1\%$ process noise, and a $6$d Lewis-Glass network model (Lewis & Glass, 1992; Gilpin, 2022), all in their chaotic regimes (see Methods A.5 for detailed parameter settings and numerical integration). From the simulated trajectories we then sampled ordinal and count observations using Eqs. (13) and (16) (with randomly drawn parameters), as well as continuous observations with $10\%$ Gaussian noise. Example reconstructions by MTF are in Fig. 2(a) and Fig. A9.
We compared the performance of MTF on these simulated data to the following setups: First, to a sequential multimodal VAE ((Kramer et al., 2022); to our knowledge the only other general approach intended for DSR from multimodal data). Second, we tried classical RNN training, where observations are provided as inputs at every time step, via truncated BPTT using modality specific decoder models, Eqs. 2-4. A third approach that can deal with multimodal observation models but, unlike TF, does not require model inversion, is ‘multiple shooting (MS)’, a method suggested in the dynamical systems literature (Voss et al., 2004), but to our knowledge not extended to multimodal data so far. We also included two other comparisons where we first transformed multimodal data to be approximately Gaussian (via Box-Cox & Gaussian kernel smoothing, see Methods A.3), and then either trained the RNN via standard BPTT-TF (Brenner et al., 2022) or by VAE-TF, similar as in MTF, but without modality-specific decoder models (labeled GVAE-TF). For comparability the same RNN architecture (the dendPLRNN, eq. 1) was used in all these comparisons.
Since in DS reconstruction we are primarily interested in invariant and long-term properties of the underlying DS (Mikhaeil et al., 2022; Platt et al., 2023; Jiang et al., 2023; Hess et al., 2023; Platt et al., 2021), for evaluation we focused on measures that capture the geometrical structure in state space ($D_{\text{stsp}}$, a type of Kullback-Leibler divergence (Koppe et al., 2019b; Brenner et al., 2022; Hess et al., 2023)) and the asymptotic temporal structure (evaluated either through the Hellinger distance, $D_{\text{H}}$, on power spectra (Mikhaeil et al., 2022; Hess et al., 2023), or auto-covariance functions for the ordinal and categorical data; Methods A.2). We also report mean-squared errors (MSE) to assess short-term ahead prediction ($10$-step-ahead $L_{2}$ prediction error (PE) for continuous and $L_{1}$ PE for ordinal data (Öğretir et al., 2022)). However, especially in chaotic systems, PEs are not suited for evaluating the system’s longer-term behavior because of the exponential divergence of nearby trajectories (Wood, 2010; Mikhaeil et al., 2022). Details of all measures are provided in Methods A.2. As evidenced in Table 1, MTF outperforms all other possible model setups, sometimes by large margins.
@@ 3.1 Evaluating MTF: Comparison to other algorithmic strategies
To investigate multimodal data integration in a more challenging setting, we next tested a scenario where continuous observations from the Lorenz-63 system were heavily distorted by Gaussian noise with $50\%$ of the data variance. At the same time, ordinal observations with $8$ variables divided into $7$ ordered categories each, $o_{nt}\in\{1\dots 7\}$, $n=1\dots 8$, were sampled using eq. 13. We then trained the dendPLRNN via MTF once with, and once without, ordinal observations. Fig. 2(b) proves that with ordinal observations on board, DS reconstruction is, in principle, possible even under these challenging conditions. The cumulative histograms of the geometric measure $D_{stsp}$ and the temporal measure $D_{H}$ in Fig. 2(c), furthermore show that inclusion of ordinal observations significantly improves reconstructions.

### v3

@@ 1 Introduction
In general, multimodal data integration is a thriving topic in many areas of artificial intelligence and ML research (Ahuja et al., 2017; Bhagwat et al., 2018; Baltrusaitis et al., 2019; Sutter et al., 2020; Shi et al., 2020; Antelmi et al., 2018; Liang et al., 2022; Lipkova et al., 2022; Xu et al., 2023; Steyaert et al., 2023; Warner et al., 2023; OpenAI, 2023). Generative models fusing multiple data channels produce a common latent code which allows for cross-modal prediction or output generation (e.g., images from language; Amrani et al. (2021); Radford et al. (2021)), improves inference by complementing information too noisy or missing in one channel through recordings from other modalities (Qian et al., 2022), or may reveal interesting links among observed modalities (Liang et al., 2015).
@@
DSR, however, goes beyond mere forecasting in that we request a dynamical model of the data-generating process which captures the temporal and geometrical structure of the true underlying system (Koppe et al., 2019b; Abarbanel, 2022; Platt et al., 2023). A successful DS reconstruction can reveal important invariant topological and geometrical characteristics of the state space the DS lives in, its long-term temporal behavior (attractors), and its sensitivity to perturbations and parameter variations. It provides a mechanistic surrogate for the observed system which can be further analyzed, simulated, perturbed, or lesioned to obtain additional insight into its inner workings (Durstewitz et al., 2023).
Often special training techniques (Mikhaeil et al., 2022; Brenner et al., 2022; Hess et al., 2023; Abarbanel, 2013, 2022; Vlachas & Koumoutsakos, 2023), regularization approaches (Schmidt et al., 2021; Platt et al., 2022; Jiang et al., 2023; Platt et al., 2023), or structural priors as in physics-informed neural networks (PINNs; Raissi et al. (2019)), are required to achieve these goals. Plain ‘out-of-the-box’ gradient descent techniques or standard variational inference will often fail to reconstruct the underlying system’s long-term properties (Arribas et al., 2020; Brenner et al., 2022; Hess et al., 2023; Platt et al., 2023; Jiang et al., 2023; Durstewitz et al., 2023), as these standard procedures do not sufficiently constrain the vector field to prevent true and reconstructed trajectories from quickly diverging. This is especially a problem in chaotic systems, where it is related to the exploding or vanishing gradient problem in RNN training (Hochreiter & Schmidhuber, 1997; Hochreiter et al., 2001; Bengio et al., 1994; Mikhaeil et al., 2022; Hess et al., 2023). Control-theoretic techniques are often used to keep model-generated trajectories on track, or to synchronize them with those produced by the observed system, whilst training (Voss et al., 2004; Abarbanel et al., 2009; Abarbanel, 2013; Verzelli et al., 2021; Platt et al., 2021; Abarbanel, 2022). Similar ideas underlie the recently proposed methods of sparse (Mikhaeil et al., 2022; Brenner et al., 2022) and generalized (Hess et al., 2023) teacher forcing (STF and GTF, respectively) for learning chaotic DS, which balance model-generated and data-inferred states during training in a way that optimally controls gradient flows and trajectory divergence (Mikhaeil et al., 2022; Hess et al., 2023; Eisenmann et al., 2023).
@@
Although important in many areas of science, an efficient framework for multimodal data integration for the purpose of DS reconstruction is currently essentially lacking. A particular challenge here is to reconcile highly efficient training techniques for DS reconstruction like STF or GTF, which usually rest on an invertible decoder model, with the need to infer a common latent code from many different types of data. Here we address this challenge through a novel formulation of the multimodal data integration problem for DS reconstruction: Building on the success of VAEs for multimodal integration (Baltrusaitis et al., 2019; Wu & Goodman, 2018; Sutter et al., 2020), we use multimodal VAEs (MVAEs) to construct a common latent representation from random variables following different distributional models. Crucially, however, rather than constructing a sequential VAE (SVAE) process that directly operates on this latent code (Kramer et al., 2022), here we employ the MVAE to create a multimodal TF signal for guiding a DSR model in training. Special loss terms and shared decoder models ensure consistency between the MVAE’s and reconstruction model’s latent codes. This not only efficiently exploits the information available in multiple different data streams for DS reconstruction (as compared to a variety of alternative techniques), but for the first time, to our knowledge, enables DS reconstructions of chaotic systems from categorical information alone.
2 Multimodal Teacher Forcing (MTF)
Techniques like STF (Brenner et al., 2022; Mikhaeil et al., 2022) or GTF (Hess et al., 2023) enable reconstructions even from challenging real-world data on which many previous methods failed. Hence, the key idea for making DS reconstruction from multiple, statistically distinct data sources work, is the generation of a multimodal TF signal for efficient training of a reconstruction model. For this we utilize an MVAE, which is then trained jointly with the DSR model through a set of shared decoder models. This general framework, illustrated in Fig. 1, is very flexible and modular. In principle it can be used with any type of DSR model for approximating the flow of the observed system, as well as with any set of decoder and encoder models.
Figure 1: MTF setup. Multimodal observations are translated via an encoder into a common latent representation, which is used for sparse TF in the DSR model’s latent space. The latent trajectory is then mapped back into observation space via modality-specific decoder models, which are shared between the MVAE and DSR model.
@@ 2 Multimodal Teacher Forcing (MTF)
While completely flexible, for our specific experiments we chose the recently introduced ‘dendritic piecewise linear RNN (dendPLRNN)’ (Brenner et al., 2022) as a reconstruction model, as it is both expressive and has a mathematically tractable, piecewise linear design (Monfared & Durstewitz, 2020; Eisenmann et al., 2023) (other types of RNNs, such as LSTMs or GRUs, may be used and give similar results, see Appx. Table B3). The dendPLRNN is defined by the $M$-dimensional latent process equation
@@
To infer this latent process equation jointly from multiple data modalities, the dendPLRNN is connected to different decoder models that take the distinct distributional properties of each modality into account. Our formulation is general, i.e. can work with any combination of continuous and/or discrete data types, and the decoders may take the form of any differentiable parametric probability model $p_{\theta}(\bullet|\bm{Z})$. For our exposition, however, we consider as concrete example time series of multivariate Gaussian ($\bm{X}$), ordinal ($\bm{O}$), [^1] Ordinal data, as often encountered in psychology or economics (Likert, 1932), are non-metric but still ordered. and count ($\bm{C}$) nature of length $T$, $\bm{Y}=\{\{\bm{x}_{1},\dots,\bm{x}_{T}\};\{\bm{o}_{1},\dots,\bm{o}_{T}\};\{\bm{c}_{1},\dots,\bm{c}_{T}\}\}$, where modalities $\bm{x}_{t}\in\mathbb{R}^{N_{1}}$, $\bm{o}_{t}\in\mathbb{Z}^{N_{2}}$, and $\bm{c}_{t}\in\mathbb{N}^{N_{3}}$ may all have different dimensions. In this case, we may take as decoders a linear Gaussian model, a cumulative link model for ordinal data, and a log-link function for Poisson data:
@@
$\bm{c}_{t}|\bm{z}_{t}$ $\sim\operatorname{Poisson}(\lambda(\bm{z}_{t})),$ (4)
The DSR model loss is now given by the sum of the negative log-likelihoods of the decoders, assuming conditional independence given latent states $\bm{z}$:
$\mathcal{L}_{\text{DSR}}=-\sum_{t=1}^{T}(\log p_{\bm{\theta}}(\bm{{x}}_{t}|\bm{z}_{1:K,t})+$ (5)
$\log p_{\bm{\theta}}(\bm{o}_{t}|\bm{z}_{1:K,t})+\log p_{\bm{\theta}}(\bm{c}_{t}|\bm{z}_{1:K,t})),$
where only the first $K$ latent states are used to allow for sharing with the MVAE as explained below, and the individual likelihood terms are specified in Appx. A.1.
@@
Training RNNs on time series is generally challenging due to the exploding and vanishing gradient problem (Bengio et al., 1994). This becomes particularly severe in DS reconstruction as for chaotic systems exploding gradients cannot be avoided even in principle, due to the positive maximum Lyapunov exponent resulting in exponential divergence of trajectories (Mikhaeil et al., 2022; Hess et al., 2023). Yet, at the same time batch length cannot be too short, as the training algorithm will then fail to capture the system’s long term behavior and limit sets (Brenner et al., 2022; Platt et al., 2022, 2023). STF (Mikhaeil et al., 2022) and GTF (Hess et al., 2023) therefore balance loss and trajectory divergence with the need to capture relevant long time scales in an optimal way. They do so by replacing forwarded-iterated latent states $\bm{z}_{t}$ by data-inferred states $\bm{\hat{z}}_{t}$ sparsely, that is only at strategically chosen time points determined from the empirically estimated maximum Lyapunov exponent in the case of STF (Mikhaeil et al., 2022), or by averaging forward-propagated and data-inferred states in an optimal way in the case of GTF (Hess et al., 2023). Data-inferred states are obtained by inverting the decoder model (e.g., $\bm{\hat{z}}_{t}=(\bm{B}^{T}\bm{B})^{-1}\bm{B}^{T}\bm{x}_{t}$ for a simple linear-Gaussian model as in Eq. 2, i.e. taking the pseudo-inverse). This forcing is then only applied during model training, not at run time (see Methods A.1 for further details).
Inverting the decoder model is in general not possible, however, in particular for discrete random variables. Moreover, in the case of multiple simultaneously observed data modalities, it is unclear how to combine the different data modalities to obtain an optimal estimate $\bm{\hat{z}}_{t}$. We therefore need another means to create a (sparse) TF signal if we would like to utilize STF or GTF for training. Here we achieve this through an MVAE for building a joint latent representation over the different data types, $\bm{\tilde{z}}_{t}\sim p(\bm{\tilde{z}}_{t}|\bm{X},\bm{O},\bm{C})$, which can be used as a TF signal at time $t$. We denote the MVAE states by $\bm{\tilde{z}}_{t}\in\mathbb{R}^{K}$ to avoid confusion with the latent dynamical process $\bm{z}_{t}\in\mathbb{R}^{M}$ generated by the reconstruction model. [^2] Note that in general we do not assume the MVAE and reconstruction model latent codes to have the same dimensionality, i.e. $K\leq M$, which, e.g., could enable the DSR model to capture additional, unobserved latent states or non-stationarity in the data; see Methods A.1 for further details. To ensure consistency between the DSR model’s and the MVAE’s latent states, the MVAE is coupled to the observations through the very same set of decoder models
@@
sharing all decoder parameters (i.e. $\bm{B},\bm{\Gamma},\bm{\beta}$ etc.) with the DSR model (see Fig. 1).
The MVAE is trained by minimizing the negative Evidence Lower Bound (ELBO)
@@
using the reparameterization trick for latent random variables (Kingma & Welling, 2014), where $\mathbb{H}_{q_{\bm{\phi}}}$ is the entropy of the approximate posterior, and assuming conditional independence of the observations given the latent states. We tested various choices for the encoder model ${q_{\bm{\phi}}}(\bm{\tilde{Z}}|\bm{Y})$, including mixture-of-experts, RNNs or transformers, but achieved best results with a temporal CNN with all data modalities concatenated (see Table B2 for results and Methods A.1 for architectural details).
@@
We thus have specified ${p_{\bm{\phi}}}(\bm{Y}|\bm{\tilde{Z}})$ through our choice of decoder models, and parameterized ${q_{\bm{\phi}}}(\bm{\tilde{Z}}|\bm{Y})$ by a temporal CNN. But what is the best choice for the prior $p_{\bm{\theta}}(\bm{\tilde{Z}})$? Crucially, to further enforce consistency between the MVAE and DSR model latent codes, $\bm{\tilde{Z}}$ and $\bm{Z}$, respectively, we assume that $p_{\bm{\theta}}(\bm{\tilde{Z}})$ is given through the DSR model (the dendPLRNN in our case). Specifically, in line with the Gaussian assumptions in the dendPLRNN, Eq. 1, we assume for the second term in Eq. 9:
$-\mathbb{E}_{q_{\bm{\phi}}}[\log p_{\bm{\theta}}(\bm{\tilde{Z}})]$ $\approx\frac{1}{L}\sum_{l=1}^{L}\sum_{t=1}^{T}\frac{1}{2}\bigl{(}\log|\bm{\Sigma}|$ (10)
@@
The total training loss is now given by the MVAE’s negative ELBO, Eq. 9, combined with the reconstruction loss (negative log-likelihood) of the DSR model (funneled through the same set of decoder models):
$\mathcal{L}_{\text{MTF}}=$ $-\mathbb{E}_{q_{\bm{\phi}}}[\log p_{\bm{\theta}}(\bm{Y}|\bm{\tilde{Z}})]-\mathbb{H}_{q_{\bm{\phi}}}(\bm{\tilde{Z}}|\bm{Y})$ (11)
$\underbrace{-\mathbb{E}_{q_{\bm{\phi}}}[\log p_{\bm{\theta}}(\bm{\tilde{Z}})]}_{\mathcal{L}_{\text{con}}}\ \ \underbrace{-\log p_{\bm{\theta}}(\bm{Y}|\bm{Z})}_{\mathcal{L}_{\text{DSR}}}$
Whilst training, we then use $\bm{\tilde{z}}_{t}$ as our TF signal for guiding the DSR model. Specifically, for STF we replace (a subset of the) latent states $\bm{z}_{t}$ by $\bm{\tilde{z}}_{t}$ at strategically chosen times $l\tau+1,\ l\in\mathbb{N}_{0}$, as illustrated in Fig. 1 (similarly, for GTF we may take a weighted average; see Hess et al. (2023) for details).
@@
Note that our usage of an MVAE here thus profoundly differs from the more conventional use in, for instance, SVAEs (Sutter et al., 2020; Kramer et al., 2022): The MVAE’s major purpose is to generate an appropriate TF control signal which guides the DSR model in training, not in modeling the process itself as in an SVAE; see Fig. A6 for an illustration of this important difference. This control of model-generated trajectories through data-inferred values has been found crucial for successful DSR (Mikhaeil et al., 2022; Brenner et al., 2022; Hess et al., 2023). This in turn necessitates the reconstruction loss $\mathcal{L}_{\text{DSR}}$, which allows for the DSR model to also make direct contact with the data, rather than only indirectly through the posterior approximation. Without it, much of the burden for recovering the dynamics would be on the (fallible) MVAE, while with it, through the sharing of decoder models, the approximate posterior will be improved through the DSR process. Ablation studies indeed further confirm that all components of the loss (Fig. A4), including $\mathcal{L}_{\text{DSR}}$, as well as an optimal choice for the TF interval (Fig. A7), are crucial to the success of the approach. Specifically, optimal TF intervals were associated with smooth and well navigable loss landscapes (Fig. A5).
3 Results
3.1 Evaluating MTF: Comparison to Other Algorithmic Strategies
We first compared MTF on moderately challenging multimodal DS reconstruction tasks [^3] Moderately challenging because many of the comparison methods actually broke down on the more difficult setups. to a variety of other possible algorithmic strategies that may be considered, all of which lack one or more key ingredients of our methodology. For this, a training and a test set of $100,000$ time steps were created from a Lorenz-63 and a Rössler system with $1\%$ process noise, and a $6$d Lewis-Glass network model (Lewis & Glass, 1992; Gilpin, 2022), all in their chaotic regimes (see Methods A.5 for detailed parameter settings and numerical integration). From the simulated trajectories, we then sampled ordinal and count observations using Eqs. 14 and 17 (with randomly drawn parameters), as well as continuous observations with $10\%$ Gaussian noise. Example reconstructions by MTF are in Fig. 2a and Fig. A10.
We compared the performance of MTF on these simulated data to the following setups: First, to a multimodal SVAE (Kramer et al. (2022); to our knowledge the only other general approach intended for DSR from multimodal data). Second, we tried classical RNN training, where observations are provided as inputs at every time step, via truncated BPTT using modality-specific decoder models, Eqs. 2-4. A third approach that can deal with multimodal observation models but, unlike TF, does not require model inversion, is ‘multiple shooting (MS)’, a method suggested in the dynamical systems literature (Voss et al., 2004), but to our knowledge not extended to multimodal data so far. We also included two other comparisons where we first transformed multimodal data to be approximately Gaussian (via Box-Cox & Gaussian kernel smoothing, see Methods A.3), and then either trained the RNN via standard BPTT-TF (Brenner et al., 2022) or by VAE-TF, similar as in MTF, but without modality-specific decoder models (labeled GVAE-TF). For comparability the same RNN architecture (the dendPLRNN, Eq. 1) was used in all these comparisons.
Since in DS reconstruction, we are primarily interested in invariant and long-term properties of the underlying DS (Mikhaeil et al., 2022; Platt et al., 2023; Jiang et al., 2023; Hess et al., 2023; Platt et al., 2021), for evaluation we focused on measures that capture the geometrical structure in state space ($D_{\text{stsp}}$, a type of Kullback-Leibler divergence; Koppe et al. (2019b); Brenner et al. (2022); Hess et al. (2023)) and the asymptotic temporal structure (evaluated either through the Hellinger distance, $D_{\text{H}}$, on power spectra (Mikhaeil et al., 2022; Hess et al., 2023), or auto-covariance functions for the ordinal and categorical data; Methods A.2). We also report mean-squared errors (MSE) to assess short-term ahead prediction ($10$-step-ahead $L_{2}$ prediction error (PE) for continuous and $L_{1}$ PE for ordinal data (Öğretir et al., 2022)). However, especially in chaotic systems, PEs are not suited for evaluating the system’s longer-term behavior because of the exponential divergence of nearby trajectories (Wood, 2010; Mikhaeil et al., 2022). Details of all measures are provided in Methods A.2. As evidenced in Table 1, MTF outperforms all other possible model setups, and in particular the multimodal SVAE by large margins, reinforcing our point about the difficulty of learning an appropriate posterior in the absence of control-theoretic guidance of the training process (cf. Fig. A6).